An investment firm recommends that a client invest in bonds rated AAA, A, and B. The average yield on AAA bonds is 4%, on A bonds 6%, and on B bonds 11%. The client wants to invest twice as much in AAA bonds as in B bonds. How much should be invested in each type of bond under the following conditions?

A. The total investment is $18,000, and the investor wants an annual return of $1,120 on the three investments.

B. The values in part A are changed to $26,000 and $1,620, respectively.

The client should invest $___ in AAA bonds, $___ in A bonds, and $___ in B bonds.
Answer #1

Let the investment in AAA be X, Let the investment in A be Y and Let the investment in B be Z.

We're given that X=2*Z (i)

A) Constraints: X+Y+Z= 18,000 => 3*Z +Y=18,000 (ii)

and 0.04*X+ 0.06*Y +0.11*Z = 1,120 => 6*Y + 19*Z= 112,000 (iii)

by (iii) - 6*(ii) we get Z= 4,000 thus X= 8,000 and Y= 6,0000.

B) By similar approach as above we get

X+Y+Z=26,000 => 3*Z+ Y=26,000 (iv)

  0.04*X+ 0.06*Y +0.11*Z = 1620=>6*Y + 19*Z= 162000 (v)

again by (v) - 6*(iv) we get Z=6,000 thus X=12000 and Y=8,000